OpenStudy (yoloshroom):

An airplane flies 490 miles with the wind and 350 miles against the wind in the same length of time. If the speed of the wind is a fixed 40 mph, what is the speed of the airplane in still air? I feel stupid.. can someone explain this to me?

OpenStudy (anonymous):

OMG! AN AMBASSADOR LIKES ANIME!

OpenStudy (compassionate):

Distance = rate x time Distance/Time = Rate "An airplane flies 490 miles with the wind and 350 miles against the wind in the same length of time. If the speed of the wind is a fixed 40 mph, what is the speed of the airplane in still air?" 490/40 - wind = 350/40 + wind Or, in other words... \[\frac{ 490 }{ 40 - r } = \frac{ 350 }{ 40 + r}\]

OpenStudy (compassionate):

Solving for r wille give you the plan in still air

Directrix (directrix):

@Compassionate I'm flying against the wind myself tonight. I don't follow how the posted equation gives the plane speed in still air. What does the variable r represent? Attached are the problem options.

OpenStudy (compassionate):

r is just the plane in still air. I jjust picked that variable.